OpenStudy (anonymous):

4(4k+2)-8(k-1)>3(2k+3)-2 a little line under the >

OpenStudy (anonymous):

sorry i edited it

OpenStudy (anonymous):

the last one someone helped me with i got wrong

OpenStudy (mertsj):

\[4(4k+2)-8(k-1)\ge 3(2k+3)-2\]

OpenStudy (anonymous):

yes thats correct

OpenStudy (mertsj):

\[16k+8-8k+8\ge6k+9-2\]

OpenStudy (mertsj):

\[8k+18\ge 6k+7\] \[2k \ge -11\]

OpenStudy (mertsj):

\[k \ge\frac{-11}{2}\]

OpenStudy (anonymous):

omg thanks so much your a life saver

OpenStudy (mertsj):

yw. And you need to learn to solve these simple problems by yourself, sweetheart.

OpenStudy (anonymous):

i got it wrong it keeps saying its wrong :(

OpenStudy (mertsj):

Check the rules for entry. Maybe it is supposed to be a decimal or a mixed number.

OpenStudy (mertsj):

No. I made a mistake.

OpenStudy (mertsj):

\[8k+16\ge6k+7\] \[2k \ge -9\]

OpenStudy (mertsj):

\[k \ge -\frac{9}{2}\]

OpenStudy (mertsj):

Was that right?